1)What is the inverse of:
y=3^x
2)What is the inverse of:
y=log(base 3)X

Answer by stanbon(48535)


To find the inverse of a function, interchange x and y
and solve for y.
1)What is the inverse of y=3^x
x = 3^y
y = log(base 3)x is the inverse
-----------------------
2)What is the inverse of:
y=log(base 3)X
x = log(base 3)y
y = 3^x is the inverse.
=================
Cheers,
Stan H.